OpenStudy (anonymous):

nPr=3024,then how to calculate the value of r?

OpenStudy (anonymous):

Same question asked here http://openstudy.com/updates/50df347fe4b0f2b98c87319f

OpenStudy (unklerhaukus):

\[^nP_r=\frac{n!}{(n-r)!}\]

OpenStudy (unklerhaukus):

Permutation

OpenStudy (anonymous):

first factorise 3024

OpenStudy (anonymous):

find the highest prime factor number in the factors here we are having 3024 = 7*3*3*3 *2*2*2*2 either start from the leading prime factor or form consecutive sequence of the factors obtained

OpenStudy (anonymous):

here we are having 3024 = 9*8*7*6

OpenStudy (anonymous):

3024 = 9*8*7*6*5! /5!= 9!/(9-4)!

OpenStudy (anonymous):

r=4